Derive an expression for density of state for particle 3-D crystal having mass m with




energy ranging between E and E + dE





1
Expert's answer
2022-02-09T16:15:22-0500

"D(E)dE=\frac{(2m)^{\frac 32}(E-E_0)^{\frac 12}}{2\pi^2 \hbar ^3}dE."